# Adaptive Prompting for Continual Relation Extraction: A Within-Task Variance Perspective

### v1

@@ 4.2 Main Results
Furthermore, we evaluate WAVE-CRE against recent successful methods in CRE, which are all rehearsal-based methods. Remarkably, without retaining training data or directly fine-tuning BERT, WAVE-CRE achieves results nearly equivalent to these state-of-the-art baselines on both datasets. Notably, in the FewRel dataset, our method even surpasses the latest rehearsal-based methods on the last task. This finding highlights the significance and effectiveness of our approach.
@@ 4.3 Detailed Analysis
To further illustrate the efficacy of our proposed task-specific prompt pools in capturing within-task variations compared to a single prompt approach for each task, we conducted experiments within the task incremental learning setting (van de Ven, Tuytelaars, and Tolias 2022). In this setting, task identities are provided, ensuring that the prompt pool is identical during training and testing. Table 2 compares the performance of models trained with and without the task-specific prompt pool. Here, “w/o Prompt Pool” represents the scenario where only a single task-specific prompt is used for each task. Our method, WAVE-CRE, trained with the task-specific prompt pool, demonstrates a substantial performance advantage over the baseline across different stages, achieving a 1.8% improvement on the final task. These results clearly highlight the effectiveness of our approach in enhancing the model’s ability to capture within-task variations.
@@
Similarly, we investigate the effect of different values of $L$ in the task incremental learning scenario. The number of prompts selected, $K$, was chosen to ensure a fair comparison by keeping the total number of experts across all experiments equal. The results are presented in Table 3. As discussed in Section 3.1, setting $L=1$ allows for flexibility in expert selection and increases the model’s expressiveness, since each expert has its own key. This is also empirically shown, as our proposal achieves the best performance among different values of $L$.
@@
We evaluate our task-ID prediction technique against the baselines EPI and HiDe-Prompt, with the results summarized in Table 4. EPI employs BERT for task identification and utilizes Mahalanobis distance to select task-specific parameters, operating under the assumption that pre-trained representations for different tasks are well-separated. However, this assumption is unrealistic, given that these representations are derived from a generic pre-trained model, which leads to poor task identity prediction accuracy. In contrast, our method incorporates a dedicated task predictor trained on synthesized query representations, resulting in significantly improved task prediction accuracy. HiDe-Prompt adopts a similar strategy to ours and demonstrates some improvement over EPI. However, as discussed in Section 3.3, its approach of grouping all relations within a task as a single class can be suboptimal. WAVE-CRE addresses this limitation by employing a task predictor with an output dimension that corresponds to the number of relations. This approach yields improvements of up to 10% on both datasets, underscoring the effectiveness of our proposed strategy.

### v3

@@ 4.2 Main Results
Furthermore, we evaluate WAVE-CRE against recent successful CRE methods, all of which are rehearsal-based. Remarkably, without retaining training data or directly fine-tuning BERT, WAVE-CRE achieves results nearly equivalent to these state-of-the-art baselines on both datasets. Notably, in the FewRel dataset, our method surpasses the latest rehearsal-based methods on the last task. This highlights the significance and effectiveness of our approach.
@@ 4.3 Detailed Analysis
To illustrate the efficacy of task-specific prompt pools in capturing within-task variations compared to a single prompt approach, we conducted experiments in the task incremental learning setting (van de Ven, Tuytelaars, and Tolias 2022). In this setting, task identities are provided, ensuring the prompt pool remains identical during training and testing. Table 2 compares models trained with and without the task-specific prompt pool. Here, “w/o Prompt Pool” represents using only a single task-specific prompt per task. Our method, WAVE-CRE, trained with the prompt pool, shows a 1.8% improvement on the final task, demonstrating its effectiveness in enhancing the model’s ability to capture within-task variations.
@@
Similarly, we investigate the effect of different values of $L$ in the task incremental learning scenario. The number of prompts selected, $K$, was chosen to ensure a fair comparison by keeping the total number of experts across experiments equal. The results are presented in Table 3. As discussed in Section 3.1, setting $L=1$ allows for flexibility in expert selection and increases the model’s expressiveness, as each expert has its own key. This is empirically shown, as our proposal achieves the best performance among different values of $L$.
@@
We evaluate our task-ID prediction technique against the baselines EPI and HiDe-Prompt, with results summarized in Table 4. EPI uses BERT for task identification and Mahalanobis distance to select task-specific parameters, assuming pre-trained representations are well-separated—an unrealistic assumption given their generic origin, leading to poor prediction accuracy. In contrast, our method trains a dedicated task predictor on synthesized query representations, significantly improving accuracy. HiDe-Prompt adopts a similar strategy and improves over EPI but groups all relations within a task as a single class, which can be suboptimal (Section 3.3). WAVE-CRE addresses this by using a task predictor with output dimensions corresponding to the number of relations, achieving up to 10% improvements on both datasets, showcasing our strategy’s effectiveness.